The length of a rectangle is 3 cm greater
than its width.The perimeter is 24 cm. Find the
dimensions of the rectangle.

Answer by Maths68(1474)

Given
Length = l =W + 3
Perimeter = 24
Width = W = ?
Perimeter of a rectangle = 2 length + 2 width
P = 2*l + 2*w
24=2*(w+3)+ 2w
24=2w+6+2w
24-6=4w
18=4w
w=18/4
w=9/2
w=4.5
l=w+3
put the value of w in above equation
l=4.5+3
l=7.5
Dimensions of rectangle
length = 7.5cm, width=4.5cm, Perimeter=24cm,
Area of rectangle =w*l=7.5*4.5=33.75cm^2